OpenStudy (anonymous):

find the equation of the normal line to f(x)=5x^2-7 at x=-1

OpenStudy (anonymous):

y=-2 right?

OpenStudy (anonymous):

idk not sure where to start

OpenStudy (anonymous):

f(x)=5x^2-7 at x=-1 =10x-0 =10x (put the value of X here which is -1) f`(x)=-10 find the value of y by puting x value in given equation =5x^2-7 =5(-1)^2-7 =5-7 y=-2 Now, Use slope equation which is y-y0=m(x-x0) y-(-2)=-10{x-(-1)} y+2=-10x-10 y=-10x-10-2 y=-10x-12 [This is ur normal line equation] To check this put the value of x in this equation y=-10x-12 y=-10(-1)-12 y=10-12 y=-2

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (anonymous):

Always Welcome....